OpenStudy (anonymous):

giving medal. Solve log 2x + log 5 = 1. Round to the nearest thousandth if necessary. a.1 b.4 c.100 d.25

OpenStudy (saifoo.khan):

Use the property: \[\log a + \log b = \log(a*b)\]